VIBRATIONAL  SPECTRA  AND  ASSIGNMENTS  OF  FUNDAMENTALS  OF  1,1-

DIFLUOROCYCLOPROPANE AND ITS d_{2} AND d_{4} ISOTOPOMERS

\underline{NORMAN C. CRAIG}, Department of Chemistry and Biochemistry, Oberlin College, Oberlin, OH 44074; DONALD C. McKEAN, School of Chemistry, University of Edinburgh, EH9 3JJ, U. K..

Infrared and Raman spectra of 1,1-difluorocyclopropane (DFCP) and its 2,2-d_{2} and d_{4} isotopomers have been recorded. Included are gas-phase Raman spectra obtained with an FT instrument and a simplified sample arrangement. A full vibrational assignment has been proposed for the three isotopomers, and normal coordinate calculations have been done in preparation for determining an equilibrium structure for this ring system. A microwave investigation provides the needed rotational constants for a full set of isotopomers of DFCP.^{a}

aA. T. Perretta, V. W. Laurie, J. Chem. Phys. 62, 2469 (1975).